# Differentially Private Zeroth-Order Methods for Scalable Large Language Model Fine-tuning

### v4

@@ 5.6. Data-free Pruning with static mask
Pruning works well in private ZO fine-tuning. We show our results in Figure 4 that under all private settings, zeroth-order fine-tuning with pruning achieves better performance (above 2%, more detailed numbers in Table 14). Especially when the privacy budget is large like $\epsilon=8,\infty$, zeroth-order fine-tuning with pruning outperforms without pruning by $4\sim 5\%$.
Figure 4. Fine-tuning RoBERTa-large model on dataset SNLI with different pruning rates. Pruning rate $r$ denotes the ratio of parameters to be tuned. Optimal $r$ scales up with the decrease of privacy budget. Detailed numbers in Table 14.
Optimal pruning rate changes with privacy budget. In our experiments, the pruning rate is seen as a hyperparameter. However, the choice of pruning rate affects the model utility on downstream tasks (see Figure 4). When the privacy budget is large (like $\epsilon=8$), useful parameters can be tuned well, small pruning rate can be chosen to lower the scale of DP noise discussed in Section 4.4. When privacy budget is small (like $\epsilon=2$), $0.5\%$ parameters can not be tuned well for large DP noise, thus bigger pruning rate $r=4\%$ should be chosen for better convergence.
We further conduct experiments on pruning-only method and rank-based important matrix with different settings of intervals (upper bound and lower bound in Algorithm 4). It is shown in Table 6 that rank-based important matrix outperforms pruning-only method under medium and large privacy budgets while behaving poorly under small privacy budgets.
[…]
@@
Figure 5. RoBERTa-large on five datasets. Dynamic pruning with a constant pruning rate fails to improve optimization.
Figure 6. RoBERTa-large on five datasets with different pruning strategies. The incremental pruning strategy is superior to other pruning strategies.
@@
5.7. Data-free Pruning with dynamic mask
@@
Dynamic pruning with constant pruning rate. We compare dynamic and constant pruning strategies under the same (constant) pruning rate. We have found that dynamic updating of the pruning mask does not necessarily lead to an improvement in accuracy with the same (fixed) pruning rate. In Table 7, we present the best accuracy results of both pruning strategies across all pruning rates. Figure 5 shows the accuracy under static and dynamic pruning strategy (constant pruning rate) varying pruning rate on five datasets.
Table 7. Experiments on RoBERTa-large (350M parameters). Dynamic pruning with a constant pruning rate fails to improve optimization.
Task SST-2 SST-5 SNLI MNLI TREC
Small Privacy budget: $\epsilon=2$
static mask 92.5 48.8 79.5 71.4 89.6
Medium Privacy budget: $\epsilon=4$
static mask 93.2 50.5 80.4 71.5 88.2
Large Privacy budget: $\epsilon=8$
static mask 93.6 49.8 81.4 73.1 90.8
We blame the failure on two main reasons: Firstly, in dynamic pruning, the parameters that were fine-tuned in the previous stage were not fully optimized and may not be selected for further fine-tuning in the next stage. As a result, the model may not be well-optimized. Secondly, the introduction of DP noise leads to the low quality of the update of the pruning mask. In Table 8, we show the proportion of the parameters fine-tuned in the current stage that are selected for fine-tuning in the next stage. Additionally, we provide the proportion of all parameters involved in the fine-tuning process compared to the total number of model parameters and the final accuracy of the model.
Table 8. RoBERTa-large on SNLI with $\epsilon=4$. “Stage 1” denotes the proportion of parameters fine-tuned in stage 1 that will be fine-tuned in the next stage (stage 2), “Total” denotes the proportion of parameters fine-tuned in the whole fine-tuning process, and “Acc” denotes the accuracy of the model. “1-2-4%$\ast$ denotes the incremental pruning strategy with increasing pruning rate 1-2-4%.
@@
Table 9. Experiments of four pruning strategies on RoBERTAa-large. All the results are the best result across all pruning rates. “static*” denotes the best result when the pruning mask is calculated before fine-tuning and remains static during the whole fine-tuning process. “dynamic(constant $r$)” and “dynamic(dynamic $r$)” denote the dynamic pruning strategy with constant and dynamic pruning rate.
@@
It is shown in Table 8 that only a minuscule proportion of parameters are selected for fine-tuning in the next stage. For example, in the first stage of 0.5-0.5-0.5%, only 0.54% of the parameters that are updated will continue to be fine-tuned in the next stage, which results in parameters not being well-tuned and they will not be fine-tuned furthermore in the following fine-tuning process. This explains why dynamic pruning with a constant rate may diminish the utility of the model compared to static pruning. However, the above issue is effectively addressed with dynamic (increasing) pruning rate and incremental strategy, which further enhance the model’s accuracy. As shown in Table 8, dynamic pruning with increasing dynamic $r$=1-2-4outperforms dynamic pruning with constant pruning rate of all three $r$=1%, 2%, and 4% (improving accuracy by 2.6%, 0.9% and 1%). Incremental dynamic pruning strategy further improves the model accuracy by 0.7%.
@@
Dynamic pruning with dynamic pruning rate and incremental strategy. We compare the best results of all four pruning strategies to explore the potential of different pruning strategies in Table 9. As shown in Figure 6, dynamic pruning with dynamic (increasing) pruning rate outperforms dynamic pruning with constant pruning rate. For example, on SNLI with $\epsilon=4$, “1-2-4%” outperforms “1-1-1%” by 2.6% for the reason that “1-2-4%” fine-tunes more parameters (1.47% compared with 3.38%) in the whole fine-tuning stage. Compared with “4-4-4%”, “1-2-4%” fine-tunes fewer parameters (5.48% compared with 3.38%) and outperforms by 1.0% for the reason that fine-tuning with a smaller pruning rate in the early stages can accelerate the convergence. We show in Figure 7 the training-set loss of “1-2-4%”, “1-1-1%”, “2-2-2%” and “4-4-4%”. “1-1-1%” converges faster in the initial stages compared to “1-2-4%”. However, since “1-2-4%” fine-tunes more parameters, it achieves better convergence with lower training-set loss.
@@
Figure 7. The training-set loss of RoBERTa-large on SNLI with privacy budget $\epsilon=4$. Dynamic pruning with increasing pruning rate “1-2-4%” performs better than both “1-1-1%” and “4-4-4%”.
Table 10. Experiments on RoBERTAa-large. 1-2-4% $\ast$ and 1-2-4% denote the incremental pruning and dynamic pruning strategy separately, both with dynamic pruning rate $r$ increasing from 1% to 2% and 4%. Incremental pruning strategy outperforms dynamic pruning strategy in most dataset under all private settings.
Task SST-2 SST-5 SNLI MNLI TREC
Small Privacy budget: $\epsilon=2$
0.5-1-2% 89.4 48.7 79.2 71.3 86.6
0.5-1-2% $\ast$ 88.5 49.0 76.7 71.6 88.4
1-2-4% 92.8 48.2 77.6 71.3 85.2
1-2-4% $\ast$ 92.2 47.7 80.1 70.7 91.2
2-4-8% 89.6 46.4 75.3 72.0 91.4
2-4-8% $\ast$ 92.8 48.6 77.5 70.6 91.4
Medium Privacy budget: $\epsilon=4$
0.5-1-2% 93.0 48.0 79.2 71.3 86.6
0.5-1-2% $\ast$ 93.6 49.1 80.0 73.1 90.4
1-2-4% 92.8 48.2 79.4 71.2 87.4
1-2-4% $\ast$ 92.9 49.3 80.1 71.6 92.0
2-4-8% 92.0 47.9 78.7 74.9 88.6
2-4-8% $\ast$ 92.8 48.6 80.6 73.9 88.6
Large Privacy budget: $\epsilon=8$
0.5-1-2% 92.2 49.8 80.8 71.9 88.0
0.5-1-2% $\ast$ 93.6 48.5 79.2 74.8 89.2
1-2-4% 92.5 48.8 79.3 72.0 89.0
1-2-4% $\ast$ 92.9 48.1 79.4 72.4 88.0
2-4-8% 92.9 46.0 81.4 74.4 89.0
2-4-8% $\ast$ 92.8 48.6 81.6 74.8 91.8
Incremental pruning strategy is a more conservative pruning strategy, it inherits the advantages of the static pruning method, fine-tuning the parameters calculated before DP fine-tuning in every stage. For example in Table 10, “2-4-8$\ast$” outperforms static pruning with $r=2\%$ by 0.2% since more parameters are fine-tuned. Compared with static pruning with $r=8\%$, “2-4-8$\ast$” achieves a $2\%$ improvement for fast convergence in the early stage. In summary, incremental pruning strategy performs better in most scenarios when it comes to private zeroth-order fine-tuning. We further present our experimental results on OPT-2.7B in Table 15.
6. Conclusion
@@ Appendix A Appendix
In Sec A.1, we provide the proof for Lemma 4.2 and the oracle complexity of both DP-ZOSO and DP-ZOPO. In Sec A.2, we will provide the experiment settings and more experiment results.
A.1. Complete Proof
In this section, we will provide detialed proof for Lemma 4.2, 4.3 and Theorem 4.4, 4.6.
Lemma A.1 (Restatation of Lemma 4.2).
@@ A.1. Complete Proof
(20) $\begin{split}\mathbb{E}[||\nabla F_{\beta}(\boldsymbol{\theta}_{t})-&\nabla\hat{f}_{\beta}(\boldsymbol{\theta}_{t},\xi_{t+1})||^{2}]\\ \leq&\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{8dC^{2}}{ePm}+\frac{64d\beta^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}.\end{split}$
Proof of lemma 4.2.
@@
(21) $\begin{split}&\mathbb{E}\left[\left\|\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1})\right\|^{2}\right]\\ =&\mathbb{E}\left[\left\|\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\frac{1}{P}\sum_{p=1}^{P}\frac{1}{m}\sum_{i=1}^{m}\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)+\mathbf{z}_{t}^{p}\right\|^{2}\right]\\ =&\frac{1}{P^{2}}\sum_{p=1}^{P}\mathbb{E}\left[\left\|\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\frac{1}{m}\sum_{i=1}^{m}\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\right]\\ &+\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}\\ =&\left\|\mathbb{E}_{\mathbf{v}}\left[\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})\right]-\frac{1}{P}\sum_{p=1}^{P}\frac{1}{m}\sum_{i=1}^{m}\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\\ &+\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}.\\ \end{split}$
@@
(22) $\begin{split}&\mathbb{E}\left[\left\|\frac{1}{2\beta}\left(f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)\right)\right\|^{2}\right]\\ =&\frac{1}{4\beta^{2}}\mathbb{E}[|f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]\\ &+\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]-f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)|^{2}]\\ \leq&\frac{1}{2\beta^{2}}\mathbb{E}\left[\left\|f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]\right\|^{2}\right]\\ &+\frac{1}{2\beta^{2}}\mathbb{E}\left[\left\|\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]-f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)\right\|^{2}\right].\end{split}$
@@
(23) $\begin{split}&\mathbb{E}\left[\left\|f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]\right\|^{2}\right]\\ =&\mathbb{E}\left[\left\|f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]\right\|^{2}\right].\end{split}$
Define $h(\boldsymbol{\theta})=f(\boldsymbol{\theta}+\beta\mathbf{v})$. Since $f$ is $L$-Lipschitz and $\mathbf{v}\in\mathcal{R}^{d}$ is sampled from standard Gaussian distribution. Then, by [(Wainwright, 2019) Proposition 3.2], we have
(24) $\mathbb{P}\left(\left|h(\boldsymbol{\theta})-\mathbb{E}[h(\boldsymbol{\theta})]\right|\geq c\right)\leq 2e^{-\frac{c^{2}}{2\beta^{2}L^{2}}}.$
@@
(26) $\begin{split}&\mathbb{E}\left[\left\|\frac{1}{2\beta}\left(f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)\right)\right\|^{2}\right]\leq 4L^{4}.\\ \end{split}$

### v5

@@ V-F Data-free Pruning with static mask
Pruning works well in private ZO fine-tuning. We show our results in Figure 6 that under all private settings, zeroth-order fine-tuning with pruning achieves better performance (above 2%, more detailed numbers in Table XVI). Especially when the privacy budget is large like $\epsilon=8,\infty$, zeroth-order fine-tuning with pruning outperforms without pruning by $4\sim 5\%$.
Figure 6: Fine-tuning RoBERTa-large model on dataset SNLI with different pruning rates. Pruning rate $r$ denotes the ratio of parameters to be tuned. Optimal $r$ scales up with the decrease of privacy budget. Detailed numbers in Table XVI.
Optimal pruning rate changes with privacy budget. In our experiments, the pruning rate is seen as a hyperparameter. However, the choice of pruning rate affects the model utility on downstream tasks (see Figure 6). When the privacy budget is large (like $\epsilon=8$), useful parameters can be tuned well, small pruning rate can be chosen to lower the scale of DP noise discussed in Section IV-D. When privacy budget is small (like $\epsilon=2$), $0.5\%$ parameters can not be tuned well for large DP noise, thus bigger pruning rate $r=4\%$ should be chosen for better convergence.
We further conduct experiments on pruning-only method and rank-based important matrix with different intervals (upper bound and lower bound in Algorithm 4). It is shown in Table V that rank-based important matrix outperforms pruning-only method under medium and large privacy budgets while behaves poorly under small privacy budgets.
TABLE V: Rank-based important matrix with different intervals on RoBERTa-large. [1.0-1.0] performs best with small privacy budget $\epsilon=2$.
@@
Figure 7: RoBERTa-large on five datasets. Dynamic pruning with a constant pruning rate fails to improve optimization.
Figure 8: RoBERTa-large fine-tuned with different pruning strategies. Incremental pruning outperforms other pruning strategies.
@@
V-G Data-free Pruning with dynamic mask
@@
Dynamic pruning with constant pruning rate. We compare dynamic and constant pruning strategies under the same (constant) pruning rate. We find that dynamic updating of the pruning mask does not necessarily lead to an improvement in accuracy with the same (fixed) pruning rate. Figure 7 shows the accuracy under static and dynamic pruning strategy (constant pruning rate) varying pruning rate on five datasets. With constant pruning rate, the static pruning strategy outperforms dynamic pruning strategy in most cases. In Table XVII, we present the best results of both pruning strategies across all pruning rates which indicates that dynamic pruning with constant pruning rate fails to improve optimization.
We blame the failure on two main reasons: Firstly, in dynamic pruning, the parameters that were fine-tuned in the previous stage were not fully optimized and may not be selected for further fine-tuning in the next stage. As a result, the model may not be well-optimized. Secondly, the introduction of DP noise leads to the low quality of the update of the pruning mask. In Table VI, we show the proportion of the parameters fine-tuned in the current stage that are selected for fine-tuning in the next stage. Additionally, we provide the proportion of all parameters involved in the fine-tuning process compared to the total number of model parameters and the final accuracy of the model.
TABLE VI: RoBERTa-large on SNLI with $\epsilon=4$. “Stage 1” denotes the proportion of parameters fine-tuned in stage 1 that will be fine-tuned in the next stage (stage 2), “Total” denotes the proportion of parameters fine-tuned in the whole fine-tuning process, and “Acc” denotes the accuracy of the model. “1-2-4%$\ast$ denotes the incremental pruning strategy with increasing pruning rate 1-2-4%.
@@
TABLE VII: Experiments of four pruning strategies on RoBERTAa-large. All the results are the best result across all pruning rates. “static*” denotes the best result when the pruning mask is calculated before fine-tuning and remains static during the whole fine-tuning process. “dynamic(constant $r$)” and “dynamic(dynamic $r$)” denote the dynamic pruning strategy with constant and dynamic pruning rate.
@@
It is shown in Table VI that only a minuscule proportion of parameters are selected for fine-tuning in the next stage. For example, in the first stage of 0.5-0.5-0.5%, only 0.54% of the parameters that are updated will continue to be fine-tuned in the next stage, which results in parameters not being well-tuned and they will not be fine-tuned furthermore in the following fine-tuning process. This explains why dynamic pruning with a constant rate may diminish the utility of the model compared to static pruning. However, the above issue is effectively addressed with dynamic (increasing) pruning rate and incremental strategy, which further enhance the model’s accuracy. As shown in Table VI, dynamic pruning with increasing dynamic 1-2-4% outperforms dynamic pruning with constant pruning rate of all three 1%, 2%, and 4% (improving accuracy by 2.6%, 0.9% and 1%). Incremental dynamic pruning strategy further improves the model accuracy by 0.7%.
@@
Dynamic pruning with dynamic pruning rate and incremental strategy. We compare the best results of all four pruning strategies to explore the potential of different pruning strategies in Table VII. As shown in Figure 8, dynamic pruning with dynamic (increasing) pruning rate outperforms dynamic pruning with constant pruning rate. For example, on SNLI with $\epsilon=4$, “1-2-4%” outperforms “1-1-1%” by 2.6% for the reason that “1-2-4%” fine-tunes more parameters (1.47% compared with 3.38%) in the whole fine-tuning stage. Compared with “4-4-4%”, “1-2-4%” fine-tunes fewer parameters (5.48% compared with 3.38%) and outperforms by 1.0% because fine-tuning with a smaller pruning rate in the early stages can accelerate the convergence. We show in Figure 9 the training-set loss of “1-2-4%”, “1-1-1%”, “2-2-2%” and “4-4-4%”. “1-1-1%” converges faster in the initial stages compared to “1-2-4%”. However, since “1-2-4%” fine-tunes more parameters, it achieves better convergence with lower training-set loss. Incremental pruning strategy is a more conservative pruning strategy, it inherits the advantages of the static pruning method, fine-tuning the parameters in the previous stage. For example in Table XIX, “2-4-8$\ast$” outperforms static pruning with $r=2\%$ by 0.2% since more parameters are fine-tuned. Compared with static pruning with $r=8\%$, “2-4-8$\ast$” achieves a $2\%$ improvement for fast convergence in the early stage. In summary, incremental pruning strategy performs better in most scenarios when it comes to private zeroth-order fine-tuning. We further present our experimental results on OPT-2.7b in Table XX.
@@
Figure 9: The training-set loss of RoBERTa-large on SNLI with privacy budget $\epsilon=4$. Dynamic pruning with increasing pruning rate “1-2-4%” performs better than both “1-1-1%” and “4-4-4%”.
Figure 10: OPT-1.3b and OPT-2.7b fine-tuned on SQuAD. DP-ZOPO performs better than other memory-efficient methods, even FT-prefix and FT under all privacy budgets.
V-H More diverse tasks
In this section, we present more experimental results pertaining to a broader range of tasks for LLMs, including question-answering, translation, and summarization tasks.
Question-answering Task: We fine-tune OPT-1.3b and OPT-2.7b on SQuAD. As shown in Figure 10 (detailed in Table XXI), model fine-tuned with DP-ZOPO has better QA accuracy than ICL, DPZero, DP-ZOSO, FT-prefix, even FT in all private settings. Our method far outperforms DPZero [1] for its large zeroth-order estimation error, constant learning rate and excessive noise injection.
TABLE VIII: OPT-1.3b and OPT-2.7b fine-tuned on Europarl.
Model OPT-1.3b OPT-2.7b
Method BLEU COMET BLEU COMET
Zero-shot 20.11 0.7562 25.36 0.7990
ICL 22.33 0.7684 27.17 0.8105
FT 21.71 0.7695 27.11 0.8135
FT-prefix 20.82 0.7609 25.21 0.8014
DPZero 20.34 0.7610 24.94 0.8040
DP-ZOSO 21.08 0.7567 25.44 0.8022
DP-ZOPO 21.19 0.7651 26.07 0.8044
TABLE IX: OPT-1.3b fine-tuned on XSum and CNN/DM. DP-ZOPO outperforms ICL and other memory efficient method FT-prefix, DPZero and DP-ZOSO.
Task Method R1 R2 RL R-SUM
XSum Zero-shot 18.66 3.51 13.95 13.93
ICL 25.15 6.92 20.07 20.07
FT 31.10 10.55 25.06 25.03
FT-prefix 27.28 9.09 22.20 22.23
DPZero 6.40 1.20 4.66 4.66
DP-ZOSO 30.13 10.34 24.43 24.45
DP-ZOPO 30.49 10.44 24.64 24.63
CNN/DM Zero-shot 32.54 15.07 23.11 23.17
ICL 34.24 16.13 24.64 24.66
FT 37.89 17.92 27.93 27.91
FT-prefix 31.29 14.89 22.59 22.55
DPZero 23.69 11.59 16.63 16.62
DP-ZOSO 31.93 15.12 23.00 23.00
DP-ZOPO 35.15 16.83 25.17 25.15
Translation Task: We fine-tune OPT-1.3b and OPT-2.7b on Europarl. As shown in Table VIII, DP-ZOPO has better results than FT-prefix, DPZero, DP-ZOSO and approaches FT. However, all DP fine-tuning (first-order and zeroth-order) methods are not as good as ICL. We attribute this situation to two reasons: 1. The OPT model is already well-trained, 2. The injection of differential privacy noise is large, although DP-ZOPO performs well among all fine-tuning methods.
Summarization Task: We fine-tune OPT-1.3b on XSum and CNN/DM with $\epsilon=4$. As shown in Table IX, models fine-tuned with DP-ZOSO and DP-ZOPO have metrics close to FT while the existing DP zeroth-order method (DPZero) fails to converge the model. DP-ZOPO outperforms ICL, FT-prefix, DPZero and DP-ZOSO. In XSum, DP-ZOPO outperforms FT-prefix by 3.21 on R1, 1.35 on R2, 2.44 on RL, 2.4 on R-SUM and approaches FT with much less GPU consumption. In CNN/DM, DP-ZOPO outperforms FT-prefix by 3.86 on R1, 1.94 on R2, 2.58 on RL and 2.6 on R-SUM.
VI Conclusion
@@ Appendix A Appendix
In Sec A-A, we provide the proof for the oracle complexity of both DP-ZOSO and DP-ZOPO. In Sec A-B, we will provide the experiment settings and more experiment results in Sec A-C.
A-A Complete Proof
In this section, we will provide detialed proof for Lemma 4, 5 and Theorem 2, 4.
Lemma 6 (Restatation of Lemma 4).
@@ A-A Complete Proof
$\begin{split}\mathbb{E}[||\nabla&F_{\beta}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta}(\boldsymbol{\theta}_{t},\xi_{t+1})||^{2}]\\ \leq&\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}+\frac{8dC^{2}}{ePm}+\frac{64d\beta^{2}L^{4}}{Pm}+\frac{\gamma^{2}}{Pm}.\end{split}$ (20)
Proof of lemma 4.
@@
$\begin{split}&\mathbb{E}\left[\left\|\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\nabla\hat{f}_{\beta_{s}}(\boldsymbol{\theta}_{t},\xi_{t+1})\right\|^{2}\right]\\ =&\footnotesize\mathbb{E}\left[\left\|\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\frac{1}{P}\sum_{p=1}^{P}\frac{1}{m}\sum_{i=1}^{m}\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)+\mathbf{z}_{t}^{p}\right\|^{2}\right]\\ =&\small\frac{1}{P^{2}}\sum_{p=1}^{P}\mathbb{E}\left[\left\|\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})-\frac{1}{m}\sum_{i=1}^{m}\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\right]\\ &+\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}\\ \small=&\small\left\|\mathbb{E}_{\mathbf{v}}\left[\nabla F_{\beta_{s}}(\boldsymbol{\theta}_{t})\right]-\frac{1}{P}\sum_{p=1}^{P}\frac{1}{m}\sum_{i=1}^{m}\mathrm{CLIP}\left(\nabla f_{\beta_{s}}^{p}(\boldsymbol{\theta}_{t},x_{t+1}^{i})\right)\right\|^{2}\\ &+\frac{dc_{2}^{2}C^{2}PT\log(P/\delta)}{\epsilon^{2}n^{2}}.\\ \end{split}$ (21)
@@
$\begin{split}&\mathbb{E}\left[\left\|\frac{1}{2\beta}\left(f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)\right)\right\|^{2}\right]\\ =&\frac{1}{4\beta^{2}}\mathbb{E}[|f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]\\ &+\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]-f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)|^{2}]\\ \leq&\frac{1}{2\beta^{2}}\mathbb{E}\left[\left\|f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]\right\|^{2}\right]\\ &+\frac{1}{2\beta^{2}}\mathbb{E}\left[\left\|\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]-f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)\right\|^{2}\right].\end{split}$ (22)
@@
$\begin{split}&\mathbb{E}\left[\left\|f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]\right\|^{2}\right]\\ =&\mathbb{E}\left[\left\|f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)-\mathbb{E}_{\mathbf{v}}\left[f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)\right]\right\|^{2}\right].\end{split}$ (23)
Define $h(\boldsymbol{\theta})=f(\boldsymbol{\theta}+\beta\mathbf{v})$. Since $f$ is $L$-Lipschitz and $\mathbf{v}\in\mathcal{R}^{d}$ is sampled from standard Gaussian distribution. Then, by [[58] Proposition 3.2], we have
$\mathbb{P}\left(\left|h(\boldsymbol{\theta})-\mathbb{E}[h(\boldsymbol{\theta})]\right|\geq c\right)\leq 2e^{-\frac{c^{2}}{2\beta^{2}L^{2}}}.$ (24)
@@
$\begin{split}&\mathbb{E}\left[\left\|\frac{1}{2\beta}\left(f(\boldsymbol{\theta}_{t-1}+\beta\mathbf{v},x)-f(\boldsymbol{\theta}_{t-1}-\beta\mathbf{v},x)\right)\right\|^{2}\right]\leq 4L^{4}.\\ \end{split}$ (26)